OpenStudy (anonymous):

What is the equation of the line that contains the points (3, –6) and (–3, 0)?

OpenStudy (anonymous):

A. y = x – 3 B. y = –x – 3 C. y = –x + 3 D. y = x + 3

OpenStudy (anonymous):

lets find the slope first slope(m) = (y2 - y1) / (x2 - x1) (3,-6) x1 = 3 and y1 = -6 (-3,0) x2 = -3 and y2 = 0 now we sub slope(m) = (0 - (-6) / (-3-3) slope(m) = (0 + 6) / -6 slope(m) = 6/-6 which is -1 now we use y = mx + b and find b slope(m) = -1 you can use either of your sets of points (3,-6) x = 3 and y = -6 now we sub again -6 = -1(3) + b -6 = -3 + b -6 + 3 = b -3 = b so your equation is : y = -x - 3 Do you have any questions ?

OpenStudy (anonymous):

yeah can you help me on this one too?? What is the equation of the line that passes through the points (0, 5) and (2, 3)?

OpenStudy (anonymous):

find the slope.. slope(m) = (y2 - y1) / (x2 - x1) (0,5) x1 = 0 and y1 = 5 (2,3) x2 = 2 and y2 = 3 now sub slope(m) = (3 - 5) / (2 - 0) can you find the slope from here ?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

check again.... slope(m) = -2/2 now what does -2/2 = ??

OpenStudy (anonymous):

1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

it equals -1. So our slope = -1. now we will use y = mx + b and we are looking for b slope(m) = -1 you can use either of your set of points (2,3) x = 2 and y = 3 now we sub y = mx + b 3 = -1(2) + b can you solve for b from here ?

OpenStudy (anonymous):

b=5

OpenStudy (anonymous):

3 = -1(2) + b 3 = -2 + b 3 + 2 = b 5 = b correct so our equation is : y = -x + 5 Do you see how I got the equation ?

OpenStudy (anonymous):

omg thanks!! can you help me with another one ??

OpenStudy (anonymous):

I suppose...what ya got ?

OpenStudy (anonymous):

What is the slope of the line represented by the equation below? y= -2/3x+3

OpenStudy (anonymous):

In y = mx + b form, the slope is in the m position. Look at your equation....y = -2/3x + 3 what number is in the m position ?

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

correct....so your slope is -2/3

OpenStudy (anonymous):

another one?

OpenStudy (anonymous):

lets see it :) We can work on it together

OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 1) and (–1, –5)?

OpenStudy (anonymous):

use the slope formula slope = (y2 - y1) / (x2 - x1) (1,1) x1 = 1 and y1 = 1 (-1,-5) x2 = -1 and y2 = -5 now sub slope = (-5 - 1) / (-1-1) can you finish and find the slope ?

OpenStudy (anonymous):

hello ??

OpenStudy (anonymous):

im doing it

OpenStudy (anonymous):

sorry...I did not mean to rush you

OpenStudy (anonymous):

3

OpenStudy (anonymous):

-6/-2 = 3 you are correct...the slope is 3

OpenStudy (anonymous):

see..you can do this...you are smart :)

OpenStudy (anonymous):

nahh i dont think so......... can you help me with another one ?

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

What is the equation of the line that has a slope of -1/3 and a y-intercept of 5/2?

OpenStudy (anonymous):

think about it....In y = mx + b form, the m is your slope and the b is your y intercept. Can you write the equation now ?

OpenStudy (anonymous):

i dont know this one

OpenStudy (anonymous):

slope = -1/3 y intercept = 5/2 y = mx + b (m is your slope and b is your y int) y = -1/3x + 5/2 Do you understand ?

OpenStudy (anonymous):

ohh i get it

OpenStudy (anonymous):

good to hear :)

OpenStudy (anonymous):

another one?

OpenStudy (anonymous):

It is not really that hard...you just need to learn the formulas. Yes..I can help with another one

OpenStudy (anonymous):

What is the equation of the line that passes through the points (1, 2/3) and (-1,-1)?

OpenStudy (anonymous):

do i need to use this formula y = mx + b?

OpenStudy (anonymous):

find the slope slope(m) = (y2 - y1) / (x2 - x1) (1,2/3) x1 = 1 and y1 = 2/3 (-1,-1) x2 = -1 and y2 = -1 time to sub slope(m) = (-1-2/3) / (-1-1) slope(m) = (-3/3 - 2/3) / (-2) slope(m) = (-5/3) / -2 slope(m) = -5/3 * - 1/2 slope(m) = 5/6 this was a little tricky because of the fraction. Look over this and see if you have any questions....then we will move one to finding the equation

OpenStudy (anonymous):

typo....then we will move on to finding the equation

OpenStudy (anonymous):

y=5/6x-1/6

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

lets see.. y = mx + b slope(m) = 5/6 use wither of your sets of points (-1,-1) x = -1 and y = -1 now we sub y = mx + b -1 = 5/6(-1) + b -1 = -5/6 + b -1 + 5/6 = b -6/6 + 5/6 = b -1/6 = b your equation is : y = 5/6x - 1/6 you got it :)

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

im done thanks

OpenStudy (anonymous):

no problem...glad I could help